A second order homogenous differential equation with constant coefficient and polynomial
coefficient is given, apply Euler-Cauchy equation to solve the given differential equation by
using the provided initial value problem (IVP).
x^2 y'' +8y' +10y =0 y(1) = 0 , y'(1) = 2.5
1
Expert's answer
2020-12-14T20:03:53-0500

"\displaystyle
x^2 y'' +8xy' +10y =0,\,\, y(1) = 0,\,\, y'(1) = 2.5\\

\textsf{Let}\,\,y = f(\ln{x}), y' = \frac{f'(\ln{x})}{x}\\

y'' = \frac{f''(\ln{x})}{x^2} - \frac{f'(\ln{x})}{x^2}\\

f"(\\ln{x}) - f'(\\ln{x}) + \\frac{8f'(\\ln{x})}{x}\\times x + 10f(\\ln{x}) = 0\\\\\n\nf"(\ln{x}) - f'(\ln{x}) + 8f'(\ln{x}) + 10f(\ln{x}) = 0\\

f"(\\ln{x}) + 7f'(\\ln{x}) + 10f(\\ln{x}) = 0\\\\\n\n\\textsf{Let}\\,\\, t = \\ln{x}\\\\\n\nf"(t) + 7f'(t) + 10f(t) = 0\\


\textsf{The auxiliary equation is}\\

m^2 + 7m + 10 = 0\\

(m + 5)(m + 2) = 0\\


m = -2, -5\\


f(t) = Ae^{-2t} + Be^{-5t}\,\, \textsf{is the general solution}\\

f(\ln{x}) = Ae^{-2\ln{x}} + Be^{-5\ln{x}}\\


y(x) = f(\ln{x}) = \frac{A}{x^2} + \frac{B}{x^5}\\


y(1) = A + B = 0\\


y'(x) = -2\frac{A}{x^3} - 5\frac{B}{x^6}\\

y'(1) = -2A - 5B = \frac{5}{2}\\


\textsf{But}\,\, A = -B\\

-2A + 5A = \frac{5}{2}\\


3A = \frac{5}{2}\\

A = \frac{5}{6}, B = -\frac{5}{6}\\


y(x) = \frac{5}{6x^2} - \frac{5}{6x^5}"